The sum of3 consecutive#s is 72 wht are the smallest of these numbers ?not sure how 2 start
Found 2 solutions by amoresroy, Alan3354:
Answer by amoresroy(361)

Let x = the smallest number
x+1 = the 2nd biggest number
x+2 = the biggest number
The sum of the 3 consecutive numbers = x + (x+1) + (x+2) = 72
3x + 3 = 72
3x = 72-3
3x = 69
x = 23
To check
23+24+25 = 72
The smallest of these numbers is 23.

Answer by Alan3354(69443)

72/3 = 24, the middle number
--> 23, 24 & 25